# Graph Anomaly Detection via Adaptive Test-time Representation Learning across Out-of-Distribution Domains

### v1

@@ 4.5 Test-Time Training
Stopping criteria. We propose an adaptive early-stopping strategy based on Maximum Mean Discrepancy (MMD) ratios between source and target domains. Our method computes MMD between target features and two distinct source feature distributions—normal and anomalous cases. The ratio of maximum to minimum MMD serves as our stopping criterion, where a higher ratio indicates better domain adaptation as it reflects stronger alignment between target samples and their corresponding source class features. The training stops when no improvement in this ratio is observed for a predefined number of epochs. Our ratio-based approach measures relative alignment between target samples and source class distributions, rather than just overall domain similarity. See Appendix A for more details about our stopping criteria.
Table 1: Statistics of the datasets.
Metric # nodes # edges # features Abnormal(%)
@@
5 Experiments
Table 2: Cross-domain (Source $\rightarrow$ Target) node-level graph anomaly detection performance of the proposed model (AdaGraph-T3) compared to state-of-the-art baselines. We have shown the results for domain pairs with both homogeneous and heterogeneous feature spaces across domains. Results show that AdaGraph-T3 significantly outperforms baselines in most settings.
Heterogeneous Features Homogeneous Features
[…]
AUROC CDA GRADE 61.80 84.96 70.38 85.53 72.06 61.27 75.51 73.80 73.16
AdaGCN 58.86 72.35 70.43 75.89 64.86 61.62 79.50 81.18 70.59
UDA-GCN 60.37 80.80 67.36 84.98 67.55 61.95 85.36 78.24 73.33
ACT 59.24 81.97 71.16 75.74 71.98 60.52 89.20 80.40 73.78
TTT TENT 59.01 76.15 71.51 83.43 66.88 58.20 75.11 79.27 71.20
GraphCL 60.37 88.46 68.71 88.81 61.91 61.64 79.05 74.20 72.89
GTrans 60.64 89.35 73.93 90.18 68.03 61.90 79.86 83.49 75.92
Ours AdaGraph-T3 62.55 91.04 80.56 94.84 86.11 60.96 93.73 90.49 82.53
AUPRC CDA GRADE 4.92 18.44 11.87 16.10 13.00 4.34 18.81 22.53 13.75
AdaGCN 4.51 9.04 12.15 5.72 10.23 4.64 22.15 35.15 12.95
UDA-GCN 4.65 6.73 10.85 22.04 10.89 5.55 26.02 27.31 14.26
ACT 4.22 8.16 13.20 26.52 12.26 4.88 33.00 28.70 16.37
TTT TENT 4.60 11.69 12.49 18.81 12.32 4.79 13.12 28.31 13.27
GraphCL 4.69 21.09 14.36 30.86 11.12 5.44 21.27 25.55 16.80
GTrans 4.78 25.95 17.02 29.11 11.91 5.15 21.03 36.20 18.89
Ours AdaGraph-T3 6.02 31.71 19.02 34.76 29.19 4.97 45.65 43.49 26.85
We will compare the proposed approach (AdaGraph-T3) against multiple baselines using six graph datasets.
Code availability: Our source code is available for reproducibility purposes as an anonymous repository. [^1] https://github.com/delaramphf/GADT3-Algorithm
5.1 Experimental Setup
Datasets. We apply six datasets from diverse domains such as online shopping reviews, including Amazon (AMZ) [25], YelpChi [27], YelpHotel (HTL), and YelpRes (RES) [8], and social networks, including Reddit (RDT) [16] and Facebook (FB)[18]. For our cross-domain analysis, we examine scenarios where feature spaces are homogeneous (same features) and heterogeneous (same features) across domains. Table 1 presents a summary of the dataset statistics. A detailed description of each dataset is provided in Appendix B.
Baselines and Evaluation Criteria. We evaluate our method against state-of-the-art baselines both for cross-domain domain adaptation and (single-graph) unsupervised anomaly detection. Cross-domain baselines include GRADE [39], AdaGCN [6], UDA-GCN [40], and ACT [36]. Test-time training baselines include TENT [34], GraphCL [43], and GTrans [13]. As GAD baselines, we consider self-supervised learning-based methods CoLA [21], SL-GAD [48], and HCM-A [11], reconstruction-based methods DOMINANT [7] and ComGA [22], and a local affinity-based method TAM [26].
We utilize two evaluation metrics to assess the performance of anomaly detection models: the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Area Under the Precision-Recall Curve (AUPRC). Higher values of AUROC and AUPRC indicate superior model performance. For all experiments, we report average results for 10 repetitions.
Implementation details. We developed AdaGraph-T3 using Python 3.11.9 and PyTorch 2.1.0. All the experiments were conducted on an NVIDIA A40 GPU with 48GB. The core model is a 2-layer GraphSAGE GNN with weight parameters optimized via the Adam optimizer [15] with a dropout rate of 0.7. The source model underwent training for 100 epochs with early stopping employed to determine the target epochs. Both the source and target models were trained with a learning rate of 0.001. We trained all baseline models using their provided source code and the hyperparameters recommended in their respective papers.
5.2 Cross-domain Graph Anomaly Detection
We present the results comparing our approach against various domain adaptation baselines in Table 2. We use different datasets as source and target domains. We incorporated a projection head into all baseline models to address feature sets’ mismatches across source and target domain datasets. This addition enables fair comparison by allowing each method to handle the discrepancy in feature spaces.
[…]
In Appendix D, we provide further insights into the performance of AdaGraph-T3. We have found that high target performance is often correlated with high source homophily. This can be explained by our use of a homophily-based self-supervised loss for test-time training.
Figure 4: Kernel PCA embeddings of node representations for (left) source domain (Amazon) and (right) target domain (Reddit). The shared representation enables the transfer of decision boundaries from the source to the target domain.
5.3 Representation Embedding
The effectiveness of our test-time training approach is visually demonstrated in Figure 4, which presents the two-dimensional Kernel PCA (polynomial kernel) embeddings of the GNN outputs for both source and target models. Importantly, we have used the same embedding space derived from the source data to project the target data, enabling the mapping of model predictions across domains.
On the left side, we observe a clear separation between normal and anomalous samples in the source domain (Amazon), indicating that the source model has effectively learned to distinguish between these classes. The right figure shows the transferability of our approach to the target domain (Reddit). Despite the inherent differences between source and target datasets, the decision boundary learned from the source domain maintains its discriminative power when applied to the target, demonstrating that our test-time training strategy successfully adapts the pre-trained source model to the new domain.
We provide additional visualizations of source and target embeddings for other datasets in Appendix E.
5.4 Early Stopping
This section evaluates our MMD ratio-based early-stopping strategy across different domain adaptation scenarios. Figure 5 illustrates how the MMD ratio evolves together with the Area Under the Precision-Recall Curve (AUPRC) during training. We marked the specific epoch when the model was selected. The results show that the ratio effectively captures the convergence of domain adaptation without the need for labeled validation data.
Table 3: Ablation study on GAD performance. Source-free AdaGraph-T3 and all baseline methods were trained and evaluated on the target dataset only (no source domain data). AdaGraph-T3 outperforms specialized GAD baselines on most datasets.
Metric (%) Method AMZ RDT FB YC Avg.
AUROC CoLA 58.98 60.28 84.34 46.36 62.49
SL-GAD 59.37 56.77 79.36 33.12 57.16
HCM-A 39.56 45.93 73.87 45.93 51.32
DOMINANT 59.96 55.55 56.77 41.33 53.40
ComGA 58.95 54.53 60.55 43.91 54.48
TAM 70.64 60.23 91.44 56.43 69.69
AdaGraph-T3 84.76 61.54 91.74 48.67 71.68
AUPRC CoLA 6.77 4.49 21.06 4.48 9.20
SL-GAD 6.34 4.06 13.16 3.50 6.76
HCM-A 5.27 2.87 7.13 2.87 4.54
DOMINANT 14.24 3.56 3.14 3.95 6.22
ComGA 11.53 3.74 3.54 4.23 5.76
TAM 26.34 4.46 22.33 7.78 15.23
AdaGraph-T3 27.53 6.01 26.13 6.14 16.45
6 Ablation Studies
We perform four ablation studies to evaluate different aspects of AdaGraph-T3. The results support the design of our model and demonstrate the relevance of each of its components.
GAD Results (Source-free). We evaluate a single-domain version of AdaGraph-T3 against six state-of-the-art GAD baselines using four datasets. Our goal is to isolate the impact of GATD3’s self-supervised learning from its cross-domain adaptation. The results are shown in Table 3. We follow the same protocol and use the same datasets as [26], so we report their results for the baselines. AdaGraph-T3 outperforms the GAD baselines across most of the datasets.
Table 4: Ablation study comparing the domain adaptation performance when the source model is trained with the supervised loss only (AdaGraph-T3-sup-source) vs. with a combination of the supervised and self-supervised loss (AdaGraph-T3). The results show that the combined loss leads to significant gains compared with the supervised loss alone.
Heterogeneous Features Homogeneous Features
[…]
AUROC GTrans 60.64 89.35 73.93 90.18 68.03 61.90 79.86 83.49 75.92
AdaGraph-T3-sup-source 60.92 90.27 75.60 92.98 70.52 59.20 92.02 84.06 78.20
AdaGraph-T3 (Ours) 62.55 91.04 80.56 94.84 81.48 60.96 93.73 90.49 81.96
AUPRC GTrans 4.78 25.95 17.02 29.11 11.91 5.15 21.03 36.20 18.89
AdaGraph-T3-sup-source 4.63 22.45 18.42 34.38 13.63 5.20 44.17 34.81 22.21
AdaGraph-T3 (Ours) 6.02 31.71 19.02 34.76 23.50 4.97 45.65 43.49 26.14
NSAW as a learnable and flexible edge weighting approach. We proposed NSAW as a learnable edge-weighting mechanism to suppress the adverse effect of non-homophily edges in anomaly detection. Table 3 compares our method against TAM [26] (using Normal Structure-preserved Graph Truncation). The results show that AdaGraph-T3 outperforms TAM using most datasets. This is evidence that NSAW is a more flexible alternative to NSGT by allowing GATD3 to learn continuous attention weights.
Class-aware regularization. In Table 5, we compare the performance of AdaGraph-T3 with and without the class-aware source regularization approach described in 4.4. The results show that the regularization helps increase the sensitivity of the source model to anomalous nodes. Regularization improves the accuracy of the model in most settings. In particular, the results show that regularization improves the performance in terms of AUPRC, which is better at capturing class imbalance than AUROC.
[…]

### v2

@@ 4.5 Test-Time Training
Stopping criteria. We propose an adaptive early-stopping strategy based on distance ratios between source and target domains. Our method computes the distance between target features and the average embeddings of two distinct source classes—normal and anomalous. The ratio of maximum to minimum distance serves as the stopping criterion, where a higher ratio indicates better domain adaptation as it reflects stronger alignment between target samples and their corresponding source class features. The training stops when no improvement in this ratio is observed for a predefined number of epochs. Our ratio-based approach measures relative alignment between target samples and source class distributions, rather than just overall domain similarity. See Appendix A.1 for details.
4.6 Theoretical Guarantee for TTT in GADT3
We propose a formal guarantee for the case of homogeneous feature spaces across domains. Under the Test-Time Training (TTT) update in GADT3, the separation margin between normal and anomalous nodes in the target graph monotonically increases. This establishes that our homophily-based self-supervised objective provably improves anomaly separability during adaptation.
We begin by introducing the notation used in the proof. Let $\theta_{s}^{*}$ denote the source encoder learned during the supervised pre-training phase, and let $\theta_{t}^{(k)}$ be the target encoder parameter learned during step $k$ of TTT. Recall that the decoder $\theta_{m}$ is kept fixed throughout TTT. For a target node $v$, we denote by $s_{t}(v;\theta_{t})$ its homophily score computed with encoder $\theta_{t}$. The anomaly separation margin at step $k$ is defined as the difference between expected values of homophily scores for normal and anomalous nodes in the graph:
$\Delta_{t}^{(k)}=\mathbb{E}_{v_{N}\in\mathcal{N}}[s_{t}(v_{N};\theta_{t}^{(k)})]-\mathbb{E}_{v_{A}\in\mathcal{A}}[s_{t}(v_{A};\theta_{t}^{(k)})],$
where $\mathcal{N}$ and $\mathcal{A}$ denote the sets of normal and anomalous nodes in the target graph, respectively. We further define the class-conditional gradient for class $C\in\{\mathcal{N},\mathcal{A}\}$ as:
$g_{C}(\theta)=\nabla_{\theta}\,\mathbb{E}_{v\in C}[s_{t}(v;\theta)].$
In particular, we denote $g_{N}^{-1}=g_{\mathcal{N}}(\theta_{s}^{*})$ and $g_{A}^{-1}=g_{\mathcal{A}}(\theta_{s}^{*})$ as the class-conditional gradients at $\theta^{\star}_{s}$. Moreover, we denote $g_{N}^{(k)}=g_{\mathcal{N}}(\theta_{t}^{(k)})$ and $g_{A}^{(k)}=g_{\mathcal{A}}(\theta_{t}^{(k)})$ the target gradients at step $k$ of TTT.
We assume that $\theta_{s}$ and $\theta_{t}$ share the same feature dimension size $(p_{s}=p_{t})$. Moreover, for the optimal source encoder $\theta_{s}^{*}$, we have the following assumptions:
Assumption 1.
(Gradient dominance at source optimum) $\|g_{N}^{-1}\|>\|g_{A}^{-1}\|$, i.e., the magnitude of the gradient induced by normal nodes is strictly greater than that induced by anomalous nodes.
Assumption 2.
(Lipschitz continuity and boundedness) The class-conditional gradients are $L$-Lipschitz continuous and uniformly bounded: $\|g_{C}(\theta_{1})-g_{C}(\theta_{2})\|\leq L\|\theta_{1}-\theta_{2}\|$ and $\|g_{C}(\theta)\|\leq G$ for all $\theta_{1},\theta_{2},\theta$ and for $C\in\{\mathcal{N},\mathcal{A}\}$.
Assumption 3.
(Initialization proximity) The target encoder initialization $\theta_{t}^{(0)}$ is sufficiently close to the optimal source encoder $\theta_{s}^{*}$.
The above assumptions are reasonable given that the source encoder $\theta_{s}^{*}$ is trained with labeled data, which naturally yields stronger gradient signals from normal nodes in the homophily-based objective. It does, however, require $\theta_{s}$ and $\theta_{t}$ to share the same feature dimension size $(p_{s}=p_{t})$. Moreover, at the source optimum, normal nodes are more homophilic and consistent, so their gradients align in direction and yield a larger $\|g_{N}^{-1}\|$ than scattered anomalous gradients.
Proposition 1 (Monotonic Margin Increase).
Under the above assumptions, for all $k\geq 0$, starting from an initial parameter $\theta_{t}^{(0)}$ sufficiently close to the optimal $\theta_{t}^{*}$, the TTT update satisfies:
$\Delta_{t}^{(k+1)}>\Delta_{t}^{(k)}.$
That is, each step of test-time training strictly increases the anomaly separation margin between normal and anomalous nodes.
5 Experiments
We evaluate the proposed approach (GADT3) against multiple baselines across six graph datasets, including methods from both cross-domain adaptation and anomaly detection. Our analysis is complemented with ablation studies on key components of the model, such as source training, NSAW, and class-aware regularization. To assess scalability, we further test GADT3 on three large-scale graphs. Additionally, we investigate the model’s sensitivity to the homophily assumption by synthetically varying homophily levels to identify the threshold at which performance begins to degrade.
5.1 Experimental Setup
Datasets. We apply six datasets from diverse domains, such as online shopping reviews, including Amazon (AMZ) (McAuley & Leskovec, 2013), YelpChi (Rayana & Akoglu, 2015), YelpHotel (HTL), and YelpRes (RES) (Ding et al., 2021), and social networks, including Reddit (RDT) (Kumar et al., 2018) and Facebook (FB)(Leskovec & Mcauley, 2012). For our cross-domain analysis, we examine scenarios where feature spaces are homogeneous (same features) and heterogeneous (different features) across domains. To demonstrate the scalability of our model, we additionally report results on three large-scale graphs: Amazon-all (AMZ-all), YelpChi-all (YC-all) (Hamilton et al., 2017), and T-Finance (TF) (Tang et al., 2022) in Table 3. A detailed description of each dataset is provided in Appendix A.2.
Baselines and Evaluation Criteria. We evaluate our method against state-of-the-art baselines both for cross-domain domain adaptation and (single-graph) unsupervised anomaly detection. Cross-domain baselines include GRADE (Wu et al., 2023), AdaGCN (Dai et al., 2022), UDA-GCN (Wu et al., 2020), and ACT (Wang et al., 2023b). Test-time training baselines include TENT (Wang et al., 2020), GraphCL (You et al., 2020), and GTrans (Jin et al., 2022). As GAD baselines, we consider self-supervised learning-based methods (CoLA (Liu et al., 2021a) and SL-GAD (Zheng et al., 2021)), reconstruction-based methods (DOMINANT (Ding et al., 2019) and ComGA (Luo et al., 2022)), and a local affinity-based method (TAM (Qiao & Pang, 2024)).
While GADT3 assumes that normal nodes are more homophilic than anomalies, most baselines (e.g., DOMINANT, CoLA, SL-GAD, GTrans) are homophily-agnostic. Yet, GADT3 outperforms them, indicating that homophily-based self-supervision provides a useful inductive bias for cross-domain anomaly detection.
We utilize two metrics to evaluate anomaly detection models: the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Area Under the Precision-Recall Curve (AUPRC). Higher values of AUROC and AUPRC indicate superior model performance. We report average results for 10 repetitions.
Implementation details. We developed GADT3 using Python 3.11.9 and PyTorch 2.1.0. All the experiments were conducted on an NVIDIA A40 GPU with 48GB. The core model is a 2-layer GraphSAGE GNN with weight parameters optimized via the Adam optimizer (Kingma & Ba, 2014) with a dropout rate of 0.7. The source model underwent training for 100 epochs with early stopping employed to determine the target epochs. Both the source and target models were trained with a learning rate of 0.001. We set $\lambda=0.001$, $\lambda_{\text{reg}}=0.1$, $\lambda_{s}=0.001$, $p=40$, and $\alpha=20$. Baselines were trained using the hyperparameters recommended in their respective papers. For hyperparameter sensitivity analysis, please refer to Appendices A.3 and A.8.
5.2 Cross-domain Graph Anomaly Detection
We present the results comparing our approach against various domain adaptation baselines in Table 1. We use different datasets as source and target domains. We incorporated a projection head into all baseline models to address mismatches in feature sets between source and target domains. This ensures a fair comparison by allowing each method to handle discrepancies in input feature spaces. In homogeneous settings, where source and target graphs share the same feature space, no projection head was used.
Table 1: Cross-domain (Source $\rightarrow$ Target) node-level graph anomaly detection performance of GADT3 compared to state-of-the-art baselines. We have shown the results for domain pairs with both homogeneous and heterogeneous feature spaces. Results show GADT3 significantly outperforms baselines in most settings.
Heterogeneous Features Homogeneous Features
[…]
AUROC
[…]
AUPRC
[…]
Our proposed method (GADT3) outperformed the baselines in most scenarios. GADT3 achieved the highest AUROC and AUPRC scores in 5 out of 6 tasks. For instance, in terms of AUROC, in the Reddit→Amazon task, GADT3 surpassed the next best method (GTrans) by 5.9%. In the Reddit→Facebook task, GADT3 outperformed the closest competitor (GTrans) by 4.7%. The only task where GADT3 did not lead in terms of AUROC and AUPRC was Facebook→Reddit, where it performed competitively but slightly below UDA-GCN. On average, our method outperformed the second-best baseline by at least 6.2 points in both AUROC and AUPRC. In Appendix A.5, we provide further insights into the performance of GADT3. We have found that high target performance is often correlated with high source homophily. This can be explained by our use of a homophily-based self-supervised loss for test-time training.
5.3 Representation Embedding
Figure 4: T-SNE embeddings of node representations for source domain (Amazon) on the left and target domain (Facebook) on the right. The source model clearly separates normal/anomalous nodes. After adaptation, the pre-trained model maintains this separation in the target domain, indicating successful transfer.
The effectiveness of our TTT approach is demonstrated in Figure 4, which presents the two-dimensional T-SNE embeddings of the GNN outputs for both source and target models. On the left side, we observe a clear separation between normal and anomalous samples in the source domain (Amazon), indicating that the source model has effectively learned to distinguish between classes. The right figure shows the transferability of our approach to the target domain (Facebook). Despite the inherent differences between source and target datasets, the adapted embeddings in the target domain preserve meaningful structure, demonstrating that our TTT strategy successfully refines the target representations using the pre-trained source model. We provide additional visualizations of source and target embeddings for other datasets in Appendix A.12.
[…]
5.4 Early-Stopping Mechanism
This section evaluates our early-stopping strategy across different domain adaptation scenarios. Figure 5 illustrates how the distance ratio evolves together with the Area Under the Precision-Recall Curve (AUPRC) during training. We marked the specific epoch when the model was selected. The results show that the ratio effectively captures the convergence of domain adaptation without the need for labeled validation data.
6 Ablation Studies
We perform four ablation studies to evaluate different aspects of GADT3. The results support the design of our model and demonstrate the relevance of each of its components.
GAD Results (Source-free). We evaluate a single-domain version of GADT3 against six state-of-the-art GAD baselines using four datasets. Our goal is to isolate the impact of GADT3’s self-supervised learning from its cross-domain adaptation. The results are shown in Table 3. We follow the same protocol and use the same datasets as Qiao & Pang (2024), so we report their results for the baselines. GADT3 outperforms baselines even in the source-free setting, as the NSAW edge weighting effectively leverages structural context to enhance anomaly separation without relying on source data.
[…]
Performance on Large-Scale Graphs. We evaluate GADT3 on three large-scale datasets and compare it against strong GAD baselines in Table 3. GADT3 outperforms all baselines on Amazon-all and T-Finance, and performs competitively on YelpChi-all.
Table 2: GADT3, trained on target only, outperforms specialized baselines on most datasets.
Metric (%) Method AMZ RDT FB YC Avg.
AUROC CoLA 58.98 60.28 84.34 46.36 62.49
SL-GAD 59.37 56.77 79.36 33.12 57.16
DOMINANT 59.96 55.55 56.77 41.33 53.40
ComGA 58.95 54.53 60.55 43.91 54.48
TAM 70.64 60.23 91.44 56.43 69.69
GADT3 84.76 61.54 91.74 48.67 71.68
AUPRC CoLA 6.77 4.49 21.06 4.48 9.20
SL-GAD 6.34 4.06 13.16 3.50 6.76
DOMINANT 14.24 3.56 3.14 3.95 6.22
ComGA 11.53 3.74 3.54 4.23 5.76
TAM 26.34 4.46 22.33 7.78 15.23
GADT3 27.53 6.01 26.13 6.14 16.45
Table 3: AUROC and AUPRC (%) on large-scale graphs. Bold indicates best.
Metric (%) Method AMZ-all YC-all TF
AUROC CoLA 26.14 48.01 48.29
SL-GAD 27.28 55.51 46.48
DOMINANT 69.37 53.90 53.80
ComGA 71.54 53.52 55.42
TAM 84.76 58.18 61.75
GADT3 86.23 55.31 62.91
AUPRC CoLA 5.16 13.61 4.10
SL-GAD 4.44 17.11 3.86
DOMINANT 10.15 16.38 4.74
ComGA 18.54 16.58 4.81
TAM 43.46 18.86 5.47
GADT3 44.98 17.77 6.23
Table 4: AUROC (%) under decreasing target-domain homophily.
[…]
0.9 62.3 81.3
0.7 62.7 79.6
0.5 60.9 79.1
0.3 58.3 63.1
0.1 58.8 63.7
Robustness Under Low Homophily Settings. Our self-supervised loss for anomaly detection is based on homophily, i.e., the tendency of similar nodes to connect. To assess the generalization of GADT3 in settings where this assumption weakens and beyond homophilic graphs, we synthetically rewired a portion of edges to reduce homophily in the AMZ and RDT datasets. As shown in Table 4, GADT3 remained robust when the target domain’s average homophily score was $\geq 0.5$, with only gradual degradation in performance below that threshold. These results show the robustness of our method while acknowledging its reliance on moderate levels of homophily.
7 Conclusion
In this paper, we have introduced GADT3, a novel TTT framework for GAD when the testing data is out-of-distribution and from heterogeneous application domains. Specifically, our approach addresses scenarios where the distributions and feature spaces differ between the source and target datasets using dataset-specific encoders. GADT3 combines the advantages of TTT with multiple innovations focused on graph anomaly detection. For instance, by leveraging the distinctive distribution pattern of node homophily between normal and anomalous nodes, we have presented a tailored SSL task along with a robust attention-based edge-weighting mechanism (NSAW) to enhance the generalization of the learned representations. Moreover, GADT3 performs model selection using a distance ratio-based early-stopping criterion. Our experiments have illustrated that GADT3 outperforms state-of-the-art graph domain adaptation baselines.
Limitations: One current limitation of our approach is its reliance on the assumption that normal nodes exhibit higher homophily compared to anomalous ones. While this pattern holds across many real-world datasets, relaxing the homophily assumption remains an important direction for future work to further enhance generalization to highly heterophilic graphs. Moreover, we will study theoretical guarantees for cross-domain GAD based on the similarity between source and target datasets in future work.
8 Acknowledgments
We gratefully acknowledge the support of the US Department of Transportation (USDOT) Tier-1 University Transportation Center (UTC) Transportation Cybersecurity Center for Advanced Research and Education (CYBER-CARE) (Grant No. 69A3552348332) and the Rice University Ken Kennedy Institute.
Appendix A Appendix
A.1 Early stopping criterion
We propose an adaptive early-stopping strategy that monitors distribution shifts between source and target domains using a distance-based metric (Dis). Let $f_{t}\in\mathbb{R}^{n_{t}\times p}$ denote the target embedding, and let $\mu_{s}^{n}\in\mathbb{R}^{p}$ and $\mu_{s}^{a}\in\mathbb{R}^{p}$ denote the pre-computed average normal and attack source embeddings, respectively, where $n_{t}$ is the number of target samples and $p$ is the feature dimension.
The distance scores are computed as:
$\text{Dis}_{n}^{i}=\|f_{t}[i]-\mu_{s}^{n}\|_{2}$ (14)
$\text{Dis}_{a}^{i}=\|f_{t}[i]-\mu_{s}^{a}\|_{2}$ (15)
The early stopping score is then defined as:
$\text{score}=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\frac{\max(\text{Dis}_{n}^{i},\text{Dis}_{a}^{i})}{\min(\text{Dis}_{n}^{i},\text{Dis}_{a}^{i})}$ (16)
For target attack samples, $\text{Dis}_{n}$ would be larger than $\text{Dis}_{a}$, making the score $\text{Dis}_{n}/\text{Dis}_{a}$, which increases as samples align better with attack source features. Conversely, for target normal samples, the score becomes $\text{Dis}_{a}/\text{Dis}_{n}$, increasing as samples align with normal source features. The training stops at epoch $T$ if:
A.2 Data description
Amazon (McAuley & Leskovec, 2013): The Amazon dataset consists of product reviews from the Musical Instruments category. Users with more than 80% helpful votes are labeled as benign entities, while those with less than 20% helpful votes are considered fraudulent entities. For each user (represented as a node in the graph), 25 handcrafted features are used as raw node features. The graph structure is defined by the U-P-U (User-Product-User) relation, which connects users who have reviewed at least one common product.
Reddit (Kumar et al., 2018): The Reddit dataset consists of forum posts from the Reddit platform, focusing on user behavior and content. In this dataset, users who have been banned from the platform are labeled as anomalies. Each post’s textual content has been vectorized to serve as one of the 64 attributes for the corresponding user or post.
Facebook (Leskovec & Mcauley, 2012): The dataset represents a social network derived from the Facebook platform. Users establish connections with other users, forming a network of relationships. Fraudulent users are assumed to be anomalies of the network. Each node represents one user with 576 node attributes.
YelpChi (Rayana & Akoglu, 2015): The dataset comprises hotel and restaurant reviews, categorized as either filtered (spam) or recommended (legitimate) by Yelp. We utilize 32 handcrafted features as raw node features for each review in the Yelp dataset. In the graph, the reviews serve as nodes. The graph’s structure is defined by the R-U-R (Review-User-Review) relation, which connects reviews posted by the same user.
YelpHotel (Ding et al., 2021): Graph dataset based on hotel reviews from Yelp, where users and hotels are nodes connected by review edges. Each review contains ratings, text, and detailed metadata about both hotels and reviewers.
YelpRes (Ding et al., 2021): Graph dataset containing restaurant reviews from Yelp, where users and restaurants are nodes connected by review edges. Each review has ratings, text, and metadata about both users and restaurants
Table 5 presents a summary of the dataset statistics.
Table 5: Statistics of the benchmark datasets used in our experiments.
  Dataset   # nodes   # edges   # features   Abnormal (%)
@@ A.2 Data description
  Amazon-all (AMZ-all)   11,944   4,398,392   25   6.87
  YelpChi-all (YC-all)   45,941   3,846,979   32   14.52
  T-Finance (TF)   39,357   21,222,543   10   4.58
A.3 Hyperparameter tuning
Figure 6: AUROC on source domains across varying $\lambda$ values. Most datasets peak at $\lambda=0.001$, indicating it as a robust default for adaptation.
We tune the main hyperparameters of GADT3, including the self-supervision weight $\lambda$, the regularization strength $\lambda_{\text{reg}}$, and the class-aware weighting factor $\alpha$. A grid search is performed using a small labeled validation set from the source (not the target) domain. Specifically, we search $\lambda\in{0.0001,0.0005,0.001,0.005,0.01}$ on a logarithmic scale, and fix $\lambda_{\text{reg}}=0.1$ and $\lambda_{s}=0.001$ based on prior ablation results. We set $\alpha$ inversely proportional to the anomaly ratio as $\alpha=1/r$, where $r$ is the proportion of anomalous nodes. As shown in Figure 6, $\lambda=0.001$ consistently achieves the best or near-best performance across source domains, making it a robust choice for cross-domain transfer.
A.4 Class-aware regularization.
In Table 6, we compare the performance of GADT3 with and without the class-aware source regularization approach described in 4.4. The results show that the regularization helps increase the sensitivity of the source model to anomalous nodes. Regularization improves the accuracy of the model in most settings. In particular, the results show that regularization improves the performance in terms of AUPRC, which is better at capturing class imbalance than AUROC.
Table 6: Ablation study of GADT3 with and without class-aware regularization. Results show AUROC (%) and AUPRC (%) across domain pairs. Bold indicates better performance.